calculate the molar enthalpy of the combustion of butane if 1.22 g of the fuel increased the temperature of a tin can calorimeter (22.5 g of tin and 200 mL of water) by 8.30 celsius
1
Expert's answer
2020-06-02T14:07:41-0400

The molar enthalpy of the combustion is the amount of energy released when 1 mole of the substance combusts in excess oxygen:

"\u2206H = -Q\/n_{C_4H_{10}}" .

The energy released can be calculated using the specific heat capacities of tin "c_{tin}" ​ (0.218 J/g°C) and of water "c_w" (4.186 J/g°C). Assuming the density of water equal 1 g/mL, the mass of water is 200 g. Therefore:

"Q = (c_{tin}m_{tin}+c_{w}m_w)\u2206T"

"Q = (0.218\u00b722.5+4.186\u00b7200)\u00b78.30 = 6.99\u00b710^3" J.

The number of the moles of butane is its mass divided by its molar mass (58.12 g/mol):

"n = m\/M = 1.22\/58.12 = 0.0210" mol.

Finally, the molar enthalpy of combustion of butane is:

"∆H = -\frac{6.99·10^3}{0.0210} = 3.33·10^5" J/mol, or 333 kJ/mol

Answer: the molar enthalpy of the combustion of butane is 333 kJ/mol.